Broussard Skateboard's sales are expected to increase by 15% from $7.0 million in 2016 to $8.05 million in 2017. Its assets totaled $5 million at the end of 2016. Broussard is already at full capacity, so its assets must grow at the same rate as projected sales. At the end of 2016, current liabilities were $1.4 million, consisting of $450,000 of accounts payable, $500,000 of notes payable, and $450,000 of accruals. The after-tax profit margin is forecasted to be 3%, and the forecasted payout ratio is 65%. Use the AFN equation to forecast Broussard's additional funds needed for the coming year. Round your answer to the nearest dollar. Do not round intermediate calculations.
Answer #1

Additional Funds Needed= (A0/S0) \Delta S - (L0/S0) \Delta S - M*(S1)*(1-Payout ratio)

A0 = Current assets
L0 = current liabilities = 1.4 million -notes payable= 900000
\Delta S = change in amount of sales
S0 = level of sales at beginning
S1 = Increased sales level
M = Profit Margin

S0= 7000000 , S1= 8050000 , \Delta S = S1 - S0 = 8050000- 7000000 = 1050000

AFN= (5000000/7000000)*1050000 - (900000/7000000)*1050000 - 0.03*(8050000)*(1-0.65)

AFN= $ 530475.00